# A time warping model for seasonal data with application to age estimation from narwhal tusks

### v1

@@ 3 Estimation
Denote the unknown parameters by $\theta:=(A,B_{0},a,b,\sigma^{2},\beta,\omega^{2},r_{1},\ldots,r_{K})$. Here we propose an algorithm for estimating $\theta$ from observations $y=(y_{0},y_{1},\ldots,y_{n})$.
@@ 3.1 Principle of the SAEM algorithm
The likelihood of the data $y_{i}$ is not explicit as model (3) depends on the hidden stochastic process $(\xi_{x})_{x\geq 0}$. A common approach for finding estimators in models with latent and unobserved data is the expectation-maximization (EM) algorithm (Dempster, Laird and Rubin, 1977). In this iterative procedure at iteration $m$, the E-step is carried out by evaluating the expected log-likelihood given data and current estimates $\hat{\theta}^{m-1}$:
$Q_{m}(\theta)$ =$\overset{def}{=}$ $Q(\theta\,|\,\hat{\theta}^{m-1})$ (8)
=$\overset{def}{=}$ $\operatorname{\mathbb{E}}_{\xi\,|\,y,\hat{\theta}^{m-1}}[\ell(y,\xi;\theta)\,|\,(y;\hat{\theta}^{m-1})]$
where $\ell$ is the complete log-likelihood of $y$ and the hidden process $\xi:=\xi_{0:n}=(\xi_{0},\ldots,\xi_{n})$. The expectation is taken wrt. the probability distribution $p(\xi|\,y,\hat{\theta}^{m-1})$. In the M-step, $Q_{m}(\theta)$ is then maximized in order to obtain new estimates $\hat{\theta}^{m}$:
@@
$=-\frac{1}{2\sigma^{2}}\sum_{i=0}^{n}(y_{i}-f(x_{i},\theta))^{2}-\frac{n+1}{2}\log\sigma^{2}+\sum_{i=1}^{n}\log p(\xi_{i}\,|\,\xi_{i-1},\theta)+\log p(\xi_{0}).$
As explained in Section 2, the transition density $p(\xi_{i}\,|\,\xi_{i-1},\theta)$ is a non-central $\chi^{2}$-distribution.
However, the conditional distribution of $\xi$ given $y$ is not explicit due to the non-linear nature of the regression function and the non-central $\chi^{2}$ distribution of the transition density. Thus, we are unable to perform the E-step in eq. (8). Instead we use (a variant of) the Stochastic Approximation EM (SAEM) algorithm (Delyon, Lavielle and Moulines, 1999; Jank, 2006). In this procedure, the E-step is replaced by two steps, a simulation (S) step and a stochastic approximation (SA) step. In the simulation step, we use a Sequential Monte Carlo (SMC) (Del Moral, Doucet and Jasra, 2006a) sampler to draw the non-observed data $\xi^{m}$, conditionally on $\hat{\theta}^{m-1}$. We then apply a stochastic approximation to equation (8):
$\tilde{Q}_{m}(\theta)=\tilde{Q}_{m-1}(\theta)+\alpha_{m}(\ell(y,\xi^{m};\hat{\theta}^{m-1})-\tilde{Q}_{m-1}(\theta))$ (10)
where $\alpha_{m}$ is a decreasing sequence of positive numbers quantifying the memory in the approximation process, fulfilling $\sum_{m}\alpha_{m}=\infty$ and $\sum_{m}\alpha_{m}^{2}<\infty$
This algorithm can be simplified when the complete likelihood belongs to the curved exponential family, which then reduces to approximating the minimal sufficient statistics of the model during the E step, while the M step is explicit through the sufficient statistics. Nevertheless, the distribution $p(\xi_{i}\,|\,\xi_{i-1})$ does not belong to a curved exponential family and we have therefore no sufficient statistics. In that case, each iteration of the EM algorithm solves
@@
However, solving the score function for the CIR process can be numerically unstable (Sørensen, 2012). We instead use martingale estimating functions (Bibby and Sørensen, 1995), leading to an approximated version of the EM algorithm.
3.2 Martingale estimating function and SAEM
At iteration $m$ of the algorithm, given the current value $\xi^{m-1}$ of the hidden process $\xi$, the M step is updated by solving the score function, i.e., by differentiating the complete log-likelihood $\ell(y,\xi^{m},\theta)$ with respect to $\theta$, the score function being approximated during the SA step.
Since the parameters $a,\beta$ and $\omega$ of the hidden process $\xi$ are difficult to estimate from the likelihood, we use a martingale estimating score function (Sørensen, 2012). The estimator is defined directly as the solution to equating the martingale estimation function to 0, mimicking the score equation. This leads to the statistics $S_{1},S_{2},S_{3}$ (Sørensen, 2012, p. 21):
@@ 3.2 Martingale estimating function and SAEM
$S_{2}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{m}+\dfrac{S_{1}}{n(1-S_{1})}(\xi_{n\Delta}^{m}-\xi_{0}^{m});$ (12)
$S_{3}=\dfrac{\sum_{i=1}^{n}(\xi_{i-1}^{m})^{-1}(\xi_{i\Delta}^{m}-\xi_{(i-1)\Delta}^{m}S_{1}-S_{2}(1-S_{1}))^{2}}{\sum_{i=1}^{n}(\xi_{(i-1)\Delta}^{m})^{-1}((\frac{1}{2}S_{2}-\xi_{(i-1)\Delta}^{m})S_{1}^{2}-(S_{2}-\xi_{(i-1)\Delta}^{m})S_{1}+\frac{1}{2}S_{2})}.$ (13)
@@
$S_{4}=\frac{1}{n}\sum_{i=0}^{n}(y_{i}-f(x_{i},\hat{\theta}^{m-1}\,|\,\xi^{m}))^{2}$ (14)
[…]
These statistics are then approximated by the stochastic approximation scheme, as in the standard SAEM. Following Ditlevsen and Samson (2014); Lavielle (2014), we choose the form
$\alpha_{m}=\begin{cases}1&m\leq m_{0}\\ (m-m_{0})^{-9/10}&m>m_{0}\end{cases}$ (15)
for the Stochastic Approximation sequence in eq. (10), where $m_{0}$ determines the iteration step wherafter memory from the previous step is included. We obtain approximations of the statistics, for $k=1,\ldots,4$:
$s_{k}^{(m)}=s_{k}^{(m-1)}+\alpha_{m}(S_{k}-s_{k}^{(m-1)}).$ (16)
@@
Since the number of cycles is proportional to $a$, we ensure that the process does not mistake random fluctuations for cycles by restricting estimates of $a$ to be between half and double of our initial estimate (using peak detection; see section 3.5). If this condition is fulfilled for the statistic associated with $a$ we update, otherwise the estimate from the previous step is kept. The update of parameters $A$, $B_{0}$ and $b$ is obtained by minimizing $S_{4}$ (substituting $\hat{\theta}^{m-1}$ by $\hat{\theta}^{m}$ and using $\hat{A}^{m-1}$, $\hat{B}_{0}^{m-1}$ and $\hat{b}^{m-1}$) using optimisation from the R-package minpack.LM (Elzhov et al., 2023).
The SAEM algorithm iterates until some stopping criteria is met. Several rules exist for determining when convergence can be assumed (Sammaknejad, Zhao and Huang, 2019). We used the average relative difference and chose the stopping time as the iteration where this average relative difference falls below a specified threshold value for 10 consecutive iterations. We chose a threshold of $\frac{1}{100}$ which ensured (empirically; found by simulations) that the majority of the trajectories of the iterative SAEM estimates had stabilized.
To summarize, our algorithm is the following. We first use SMC to filter multiple trajectories of $\xi$ called particles, using the exact transition density. This algorithm is detailed in Section 3.3. Then, using the particles $\xi^{m}$, we compute $r_{k}=\tilde{y}(x^{*}_{k})-f_{0}(x^{*}_{k};\theta^{m-1})$ where $\tilde{y}(x^{*}_{k})$ is our estimate of the maximum amplitude at peak $x^{*}_{k}$ and $x^{*}_{k}$ is derived from the set (6) using $\xi^{m}$. This computation is also detailed in Section 3.3. Once the latent variables are filtered, we proceed to the stochastic approximation step. In the final maximization (M) step, we update the estimates $\hat{\theta}^{m}$. This procedure reiterates until some stopping criteria is met. The basic estimation procedure, tracing the steps above, is summarized in Algorithm 1. Later we generate confidence intervals for the parameters and other derived quantities. This subsequent step is summarized in Algorithm 4.
@@
Result: ${\xi},\hat{\theta}:=(\hat{A},\hat{B}_{0},\hat{b},\hat{a},\hat{\sigma}^{2},\hat{\rho},\hat{\omega}^{2},\hat{r}_{1},\ldots,\hat{r}_{K})$
Step(1)  Initialize parameters $\hat{\theta}^{0}$ and the unknown process ${\xi}^{0}$ (see Section 3.5);
@@
       Step(2)   Update $\xi^{m}$ given $\hat{\theta}^{m-1},\hat{r}^{m-1}_{k}$ using SMC (Algorithm 2) ;
       Step(3)   Update $\hat{r}_{k}^{m}$ given $\xi^{m}$, $\hat{\theta}^{m-1}$ using Algorithm 3;
       Step(4)   Do the stochastic approximation of statistics $S_{1},\ldots,S_{4}$;
       Step(5)   Update $\hat{a}^{m},\hat{\rho}^{m},(\hat{\omega}^{2})^{m},(\hat{\sigma}^{2})^{m}$ using the statistics $s_{1}^{(m)}\ldots,s_{4}^{(m)}$;
       Step(6)   Update $\hat{b}^{m},\hat{A}^{m},\hat{B}_{0}^{m}$ by optimizing $S_{4}$ ; $m\leftarrow m+1$

### v2

@@ 3 Estimation
The estimation method has three steps, a preprocessing step, an initialization step (section 3.2) and the main algorithm (section 3.3, 3.4 and 3.5) We finally consider how to validate the model and quantify uncertainty using bootstrapping (section 3.6).
3.1 Preprocessing
Trend removal: If the data show clear trends, these should be removed before the analysis. For our data, we only centered the signal around $y=0$ by subtracting the average $\bar{y}$. Amplitude normalization: To obtain approximately constant amplitudes, we assume that the signal (roughly) concentrates around a narrow band of frequencies and that the amplitude process is slowly evolving, in order to use the Hilbert transform to compute the envelope of the signal (Osgood, 2007, p. 318–320). We then apply a rolling average with a window size of approximately $10\%$ of the data to smooth the upper envelope, which we denote $z_{i}=z(x_{i})$. The signal used for further analysis is then normalized as $y_{i}\leftarrow(y_{i}-\bar{y})/z_{i}$. This implies that $\max(f(x))=A+B\approx 1$, and we therefore assume that B=1AB=1-A. After the analysis, the fitted curve can be transformed back to obtain a fit to the original data.
3.2 Parameter initialization
The loss function will likely exhibit many local maxima due to the trigonometric functions, and proper initialization will therefore assist the algorithm in convergence towards a global optimum. Obtaining good initial estimates is a non-trivial task, and typically involves heuristic approaches. Here we present our chosen initialization, denoted $\hat{\theta}^{0}$. In steps (1) and (2) we use a smoothed signal, denoted $\tilde{y}$, by performing loess regression with a span equal to 2% of the data. This is implemented with the loess function in base R.
(1) Amplitudes $A,B$ and phase offset $b$. The signal (3) attains maximum and minimum values $f_{max}:=A+B$ and $f_{min}:=-B-A^{2}/(8B)$ assuming, in case of the minimum, that $A\leq 4B$ (see supplementary material). Values $f_{max}$ and $f_{min}$ are unknown, but can be approximated by $\max(\tilde{y})=\tilde{y}_{max}$ and $\min(\tilde{y})=\tilde{y}_{min}$. Solving for $A$ and $B$, we obtain $\hat{B}=\frac{1}{9}\left(\tilde{y}_{max}-4\tilde{y}_{min}\pm\sqrt{(\tilde{y}_{max}-4\tilde{y}_{min})^{2}-9\tilde{y}_{max}^{2}}\right)$ and $\hat{A}=\tilde{y}_{max}-\hat{B}$, which offer two possible solutions, $(\hat{A}_{1},\hat{B}_{1})$ and $(\hat{A}_{2},\hat{B}_{2})$. Define the set of amplitude candidates
$S=\{(A,B)\in((\hat{A}_{1},\hat{B}_{1}),(\hat{A}_{2},\hat{B}_{2}),(\tilde{y}_{max},0))\,:\,A>B\},$ (6)
where we added $(\tilde{y}_{max},0)$ to ensure $S$ is non-empty. We initialize $\hat{A}^{0},\hat{B}^{0}$ by drawing a random pair from $S$. The phase offset is initialized to $\hat{b}^{0}=\pi$ to ensure that the initial grid of $b$ candidates span the entire unit circle in the first run of the algorithm, see Section 3.5.
(2) Measurement error: $\sigma^{2}$. The variance of the measurement noise is approximated from the smoothed signal $\tilde{y}$, setting $(\hat{\sigma}^{2})^{0}=\frac{1}{n}\sum_{i=1}^{n}(y(x_{i})-\tilde{y}(x_{i}))^{2}$.
(3) Hidden process parameters: $a,\beta,\omega^{2}$. With initial estimates for all other parameters, we now run SMC (Algorithm 2) with a fixed number of particles $n_{p}$, but where each particle is initialized with a random pair $(a,\beta,\omega^{2})\in(\tfrac{2\pi C_{min}}{n\Delta},\tfrac{2\pi C_{max}}{n\Delta})\times(0.01,0.5)\times(0.01,0.3)$ obtained from rejection sampling under the Feller constraint $2a\beta>\omega^{2}$. Here, $C_{min/max}$ is our prior belief on the lower/upper number of cycles in the signal. We resample among the particles at each step and select a final candidate based on the final weights. The parameters associated with the selected particle define our initial estimates $(\hat{a}^{0},\hat{\beta}^{0},(\hat{\omega}^{2})^{0})$.
@@ 3.3 Principle of the SAEM algorithm
Here, we propose an algorithm for estimating $\theta=(A,B,a,b,\sigma^{2},\beta,\omega^{2})$ from observations $y=(y_{0},y_{1},\ldots,y_{n})$.
The likelihood is not explicit as model (3) depends on the hidden stochastic process $(\xi_{x})_{x\geq 0}$. A common approach for estimating in models with unobserved variables is the expectation-maximization (EM) algorithm (Dempster, Laird and Rubin, 1977). In this iterative procedure at iteration $m$, the E step is carried out by evaluating the expected log-likelihood given data and current estimates $\hat{\theta}^{m-1}$:
$Q_{m}(\theta)$ $\overset{def}{=}$ $Q(\theta\,|\,\hat{\theta}^{m-1})$ (7)
$\overset{def}{=}$ $\operatorname{\mathbb{E}}_{\xi\,|\,y,\hat{\theta}^{m-1}}[\ell(y,\xi;\theta)\,|\,(y;\hat{\theta}^{m-1})],$
where $\ell$ is the complete log-likelihood of $y$ and the hidden process $\xi=(\xi_{0},\ldots,\xi_{n})$. The expectation is taken with respect to the probability distribution $p(\xi|\,y,\hat{\theta}^{m-1})$. In the M-step, $Q_{m}(\theta)$ is then maximized in order to obtain new estimates $\hat{\theta}^{m}$:
@@
$=-\frac{1}{2\sigma^{2}}\sum_{i=0}^{n}(y_{i}-f(x_{i},\theta))^{2}-\frac{n+1}{2}\log\sigma^{2}+\sum_{i=1}^{n}\log p(\xi_{i}\,|\,\xi_{i-1},\theta)+\log p(\xi_{0}).$ (9)
The transition density $p(\xi_{i}\,|\,\xi_{i-1},\theta)$ is a non-central $\chi^{2}$-distribution, see Section 2.
The conditional distribution of $\xi$ given $y$ is not explicit due to the nonlinear nature of the regression function and the non-central $\chi^{2}$ distribution of the transition density. Thus, we cannot perform the E step in Eq. (7). Instead, we use a variant of the Stochastic Approximation EM (SAEM) algorithm (Delyon, Lavielle and Moulines, 1999; Jank, 2006), the novelty being the introduction of a martingale estimating function to solve the M step. In this procedure, the E step is replaced by two steps, a simulation (S) step and a stochastic approximation (SA) step. In the simulation step, we use a Sequential Monte Carlo (SMC) (Del Moral, Doucet and Jasra, 2006a; Gordon, Salmond and Smith, 1993) sampler to draw the non-observed data $\xi^{m}$, conditionally on $\hat{\theta}^{m-1}$. We then apply a stochastic approximation to Eq. (7):
$\tilde{Q}_{m}(\theta)=\tilde{Q}_{m-1}(\theta)+\alpha_{m}(\ell(y,\xi^{m};\hat{\theta}^{m-1})-\tilde{Q}_{m-1}(\theta)),$ (10)
where $\alpha_{m}$ is a decreasing sequence of positive numbers quantifying the memory in the approximation process, fulfilling $\sum_{m}\alpha_{m}=\infty$ and $\sum_{m}\alpha_{m}^{2}<\infty$, which ensures convergence of the algorithm (Delyon, Lavielle and Moulines, 1999).
The algorithm can be simplified when the complete likelihood belongs to a curved exponential family, which then reduces to approximating the minimal sufficient statistics of the model during the E step, while the M step is explicit through the sufficient statistics. Nevertheless, the distribution $p(\xi_{i}\,|\,\xi_{i-1})$ does not belong to a curved exponential family, and we do not have sufficient statistics. In that case, each iteration of the EM algorithm solves
@@
However, solving the score function for the square-root process can be numerically unstable, because it involves a Bessel function of the first kind (Ditlevsen and Lansky, 2006). The square-root diffusion belongs to the class of Pearson diffusions, which are statistically tractable and well-behaved (Forman and Sorensen, 2008). Martingale estimating functions provide a useful alternative with well established asymptotic properties (consistency, asymptotic normality) similar to the maximum likelihood estimator (Sørensen, 2012). Furthermore, estimating functions have shown useful for parameter estimation in the square-root process when only observed through the integrated process (Ditlevsen and Sørensen, 2004), which is the case of the growth process $g$. Therefore, we propose to use a martingale estimating function (Bibby and Sørensen, 1995; Sørensen, 2012) to solve the M step, leading to an approximate version of the EM algorithm.
3.4 Martingale estimating function and SAEM-variant
At iteration $m$ of the algorithm, given the current estimate $\xi^{m-1}$ of the hidden process $\xi$ obtained using SMC (Section 3.5), the M step is updated by solving the score function, i.e., by differentiating the complete log-likelihood $\ell(y,\xi^{m},\theta)$ with respect to $\theta$, the score function being approximated during the SA step.
The estimators of parameters $a,\beta$ and $\omega$ of the hidden process $\xi$ are defined directly as the solution to the martingale estimation function, mimicking the score equation. This leads to the statistics $S_{1},S_{2},S_{3}$ (Sørensen, 2012, p. 21):
@@ 3.4 Martingale estimating function and SAEM-variant
$S_{2}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i\Delta}^{m}+\dfrac{S_{1}}{n(1-S_{1})}(\xi_{n\Delta}^{m}-\xi_{0}^{m});$ (12)
$S_{3}=\dfrac{\sum_{i=1}^{n}(\xi_{(i-1)\Delta}^{m})^{-1}(\xi_{i\Delta}^{m}-\xi_{(i-1)\Delta}^{m}S_{1}-S_{2}(1-S_{1}))^{2}}{\sum_{i=1}^{n}(\xi_{(i-1)\Delta}^{m})^{-1}((\frac{1}{2}S_{2}-\xi_{(i-1)\Delta}^{m})S_{1}^{2}-(S_{2}-\xi_{(i-1)\Delta}^{m})S_{1}+\frac{1}{2}S_{2})}.$ (13)
@@
$S_{4}=\frac{1}{n}\sum_{i=0}^{n}(y_{i}-f(x_{i},\hat{\theta}^{m-1}\,|\,\xi^{m}))^{2},$ (14)
where $g^{m}(x_{i})=\int_{0}^{x_{i}}\xi^{m}_{s}ds$. These statistics are then approximated by the stochastic approximation scheme, as in the standard SAEM. Following Ditlevsen and Samson (2014); Lavielle (2014), we choose the form
$\alpha_{m}=\begin{cases}1&m\leq m_{0}\\ (m-m_{0})^{-0.8}&m>m_{0}\end{cases},$ (15)
for the Stochastic Approximation sequence in Eq. (10), where $m_{0}$ determines the first iteration that includes memory from the previous step. We obtain approximations of the statistics,
$s_{k}^{(m)}=s_{k}^{(m-1)}+\alpha_{m}(S_{k}-s_{k}^{(m-1)}),\quad k=1,\ldots,4.$ (16)
@@
The number of cycles is proportional to $a$. We ensure that the process does not mistake random fluctuations for cycles by restricting estimates of $a$ to be between half and double of the initial estimate (see Section 3.2). If this condition is satisfied for the statistic associated with $a$, we update, otherwise the estimate from the previous step is kept. This can, of course, be adapted to the requirements of specific applications.
To update $A$ and $B$, we use that after normalization B=1AB=1-A. Substituting in Eq. (3) we obtain
$f(x_{i},\theta)=Aw_{i}+v_{i},$ (17)
where $w_{i}=\sin(g(x_{i})+b)+\cos(2g(x_{i})+2b)$ and $v_{i}=-\cos(2g(x_{i})+2b)$. Eq. (17) allows for a linear regression with constraint $A\in(0.5,1)$ since $A+B=1$ and $A>B>0$. Then
$\hat{A}^{m}$ $=\operatorname*{arg\,min}_{A\in\left(0.5,1\right)}\sum_{i=1}^{n}\left(y_{i}-Aw_{i}-v_{i}\right)^{2}$ (18)
$=\min\left(1,\;\max\left(0.5,\;\frac{\sum_{i=1}^{n}w_{i}(y_{i}-v_{i})}{\sum_{i=1}^{n}w_{i}^{2}}\right)\right),$ (19)
where $\sum_{i=1}^{n}w_{i}(y_{i}-v_{i})/\sum_{i=1}^{n}w_{i}^{2}$ is the least squares estimator in the unconstrained problem. Finally, $\hat{B}^{m}=1-\hat{A}^{m}$.
It turns out that the initial phase $b$ and the growth process $g(x)$, and consequently $\xi$, are not individually identifiable and must be estimated jointly. In section 3.5, we go into detail on how we can modify the SMC sampler to simultaneously sample $\hat{b}^{m}$ in iteration $m$, by making a grid of plausible $b$ candidates, and run a single SMC filter for each $b$. We call this SMC+. A brief example showing why sequential estimation of $g(x)$ and $b$ is not feasible is given in the supplementary material.
The SAEM algorithm iterates until some stopping criteria is met. Several rules exist for determining when convergence can be assumed (Sammaknejad, Zhao and Huang, 2019; Dhull and Kumar, 2022; Lixoft SAS, 2024; Delyon, Lavielle and Moulines, 1999). One common rule is to ensure that the maximum of the relative differences of all parameters between consecutive iterations falls below some threshold to ensure that the SAEM estimates have stabilized. We took the threshold to be $10^{-4}$ following Searle, Casella and McCulloch (1992). We also make a simulation study to investigate the sensitivity to this threshold choice, see supplementary material.
In summary, the algorithm is as follows. We first use SMC to filter multiple trajectories of $\xi$ called particles, using the transition density. This is repeated over a grid of candidates for the phase parameter $b$. This algorithm is detailed in section 3.5. Once the latent process is filtered, we proceed to the stochastic approximation step. In the final maximization (M) step, we update $\hat{\theta}^{m}$. This procedure is repeated until some stopping criteria are met. The basic estimation procedure is summarized in Algorithm 1. Later we construct confidence intervals for the parameters and other derived quantities. This is summarized in Algorithm 4.
@@
Result: ${\xi},\hat{\theta}=(\hat{A},\hat{B},\hat{b},\hat{a},\hat{\sigma}^{2},\hat{\rho},\hat{\omega}^{2})$
(1)    Initialize parameters $\hat{\theta}^{0}$ and the unknown process ${\xi}^{0}$ (Section 3.2);
@@
 (2)   Update $\hat{b}^{m},\xi^{m}$ given $\hat{\theta}^{m-1}$ using SMC+ (Algorithm 2) ;
 (3)   Do the stochastic approximation of statistics $S_{1},\ldots,S_{4}$;
 (4)   Update $\hat{a}^{m},\hat{\rho}^{m},(\hat{\omega}^{2})^{m},(\hat{\sigma}^{2})^{m}$ using the statistics $s_{1}^{(m)}\ldots,s_{4}^{(m)}$;
 (5)   Update $\hat{A}^{m}$ using Eq. (19) and set $\hat{B}^{m}=1-\hat{A}^{m}$ ;
 $m\leftarrow m+1$